OpenStudy (anonymous):

What is the distance between points (4, −7) and (−5, 9)?

OpenStudy (anonymous):

\[\sqrt{(4+5)^2 +(-7-9)^2}\]

OpenStudy (anonymous):

\[\sqrt{81+256}\]

OpenStudy (anonymous):

\[\sqrt{337}\]